# Advances and Open Challenges in Federated Learning with Foundation Models

### v1

@@ VI-D2 Contribution Evaluation for Centralized FMs
Automated Evaluation: Automated evaluation of FMs commonly adopts standard metrics or indicators and evaluation tools to assess model performance. These include accuracy, fairness, ROUGE [189] and BERTScore [190]. For instance, the BLEU score [191] has been leveraged to measure the similarity and quality between the text generated by the FM and the reference text in a machine translation task. Due to its simplicity and automatic computing, this evaluation metric is widely adopted by existing FM evaluation effort. In addition, with the development of FMs, more advanced automated evaluation techniques are emerging. For example, Lin et al. [192] proposed a unified multidimensional automated evaluation approach, LLM-EVAL, for domain conversations. PandaLM [193] trained an additional LLM to evaluate different models, which is suitable for reproducible and automated language model evaluation.
Human Evaluation: Human evaluation aims to assess the quality and accuracy of FM-generated results through human participation. Different from automated evaluation, human evaluation focuses more on specific application scenarios and can provide more comprehensive results. Usually, evaluators consist of experts, researchers or the target users. Recent human evaluation methods involve tasks including generation, summarization and analogical reasoning tasks. Bubeck et al. [194] conducted various human-crafted tests on GPT-4. They observed that GPT-4 performs close to or even exceeds human performance on various tasks. In practice, automated evaluation and human evaluation are considered for adoption on a case-by-case basis.
@@
During fine-tuning, the training data might not be equally important for a pre-trained FM. On the one hand, pre-trained FMs are prone to significant performance degradation in the presence of noisy data. This effect can be further amplified when noisy or harmful samples are highly influential to the model. On the other hand, specific knowledge embedded within some samples might have been extracted after some training rounds. Thus, they can be ignored afterwards without affecting final fine-tuned model performance. Therefore, it is important to identify noisy samples and important samples in the fine-tuning data to improve training efficiency and model performance. Jain et al. [195] evaluate FMs from the data-based perspective by eliminating the need for laborious labeling of new data. In [196], an efficient SV-based data evaluation method has been proposed. It achieves this design goal through an efficient sampling-based method that aggregates SVs computed from subsets for valuation of the entire training set, and a value transfer method that leverages value information extracted from a simple classifier trained using representations from the target language model. DataInf [179] is a computationally efficient influence approximation method that is based on an easy-to-compute closed-form expression. It can be easily applied to FMs.
@@ VI-D3 Contribution Evaluation for FedFM
As for evaluating FedFM, so far, there have been few relevant works. The work [197] proposed a logic rule learning approach to select the optimal chain-of-thoughts prompts for improving the interpretability of federated prompt selection for multi-domain FMs. They cast this problem as a bi-level program, and solve it through variational expectation maximization. The work [198] proposed to evaluate FedFM by comparing the performance of the federated parameter-efficient fine-tuned model with traditional centralized fine-tuning methods. FedIT [199] leverages FL as the learning framework for FM instruction tuning and conducted studies on the widely used GPT-4 to exploit the heterogeneous and diverse sets of instructions. FedNLP [200] is a benchmark framework for evaluating FL methods on four common formulations of NLP tasks: text classification, sequence tagging, question answering and seq2seq generation. They proposed a universal interface between Transformer-based language models (e.g., BERT, BART) and FL methods under various non-IID partitioning strategies.
@@ VIII-B Communication Challenges facing FedFM
Increased Communication Overhead vs. Bandwidth Constraints: Bandwidth constraints present a critical challenge to FedFM [226]. The sheer size of FMs necessitates large transmission overhead, which is compounded by the need for frequent communication across distributed training devices (i.e., clients and servers). While strategies like model compression (using techniques like pruning, quantization, and knowledge distillation) aim to reduce the size of the models, and thus the network load [227], they only partially alleviate the bandwidth demand. Sparse communication techniques [228], which focus on sharing only essential model updates, help to some extent but do not entirely solve the problem. Asynchronous communication strategies [229] and edge computing [230] can reduce network congestion and data transmission volume. Yet, they cannot fully compensate for the inherent high transmission requirements of FedFM. Efficient coding schemes optimize data transmission [231], but the fundamental challenge of transmitting large volumes of data in a bandwidth-limited environment persists. Collectively, these approaches make strides in addressing bandwidth issues, but they do not completely solve the challenge, highlighting the need for innovative solutions in the field of FedFM [15].
@@
Privacy Issues: Privacy concerns in FedFM remain a significant challenge, despite the ongoing privacy protection research in traditional FL settings. While FL enhances privacy by sharing model updates rather than raw data, there are still vulnerabilities during interactions between FL participants [242]. For instance, adversaries can exploit these vulnerabilities to perform attacks such as membership inference or gradient leakage [242, 243, 244], aiming to extract local training data from devices. Existing countermeasures, such as secure multi-party computation (SMC) and homomorphic encryption (HE), while effective in enhancing FL privacy to some extent, fall short in fully addressing these malicious attacks. As detailed in [245], HE can protect against data leakage but is less effective against more sophisticated threats like membership inference and gradient leakage attacks. However, it cannot be efficiently applied on large-scale FMs. Despite these efforts in improving privacy protection, the evolving nature of threats and the complexity of FL systems, particularly in the context of FMs, means that privacy concerns remain a pressing and unresolved issue in FedFM. The development of more advanced and resilient privacy-preserving mechanisms continues to be a critical need in this field.
@@ VIII-C Promising Direction: Quantum Computing for FedFM
The integration of quantum computing with FedFM presents a transformative opportunity to enhance FedFM [246]. This integration could revolutionize the efficiency and capability of these models [247]. In this section, we envision how four promising quantum computing techniques (Table X): 1) quantum machine learning (ML), 2) quantum optimization, 3) quantum security and 4) a combination of quantum entanglement and quantum aggregation hold promise to improve the efficiency, trustworthiness and incentives aspects of FedFM.
@@
Quantum Entanglement and Quantum Aggregation for FedFM: Quantum entanglement offers a novel approach to synchronizing updates in FedFM. This quantum phenomenon, where two or more particles become interconnected so that the state of one instantly influences the others, can be useful for FedFM in multiple ways. Theoretically, quantum entanglement might enable immediate synchronization of model updates across different nodes in a federated network [263]. Using entangled particles to represent model parameters allows changes in one node to be instantly reflected in all other entangled nodes, ensuring real-time synchronization of updates. Reaching consensus on model updates among distributed nodes is a significant challenge in FedFM. Quantum entanglement offers an efficient mechanism for achieving instant consensus. Quantum entanglement can achieve highly resource-efficient synchronization by reducing communication overhead and allowing for more effective error correction [264]. This can save bandwidth and computational resources, with quantum entanglement-assisted aggregation potentially correlating model updates from different nodes more efficiently. While still in the theoretical stage, the potential impact of quantum entanglement on FedFM could revolutionize the way model updates are synchronized, with highly enhanced efficiency, security and resource management.
Quantum aggregation in FedFM involves leveraging quantum computing principles for aggregating model updates from various nodes. This process aims to compute a global model update that effectively integrates local model changes from each node, potentially enhancing both efficiency and security. Key approaches in quantum aggregation include the follows. Quantum superposition [265] and quantum interference [266] involve using quantum superposition to represent multiple model updates on a single quantum state. This state can then be processed to compute a global model update. Quantum interference, where quantum states interfere constructively or destructively, can facilitate efficient FedFM aggregation. Quantum summation algorithms can perform aggregation tasks more efficiently than classical algorithms [267]. Leveraging quantum cryptography techniques, quantum secure aggregation of model updates can preserve privacy while enabling effective aggregation. While detailed studies specific to FedFM and quantum aggregation are not yet available, the principles of quantum computing offer promising pathways for enhancing the efficiency and security of FedFM model aggregation.
@@ VIII-D Promising Direction: Evaluation of FedFM
Evaluation under Complex Threat Models: Most of the current FL evaluation approaches are built on the simple threat model of semi-honest participants. This renders them susceptible to scenarios in which the server or clients are malicious or colluding. It is imperative to relax this simplifying assumption to enable future FedFM evaluation approaches to handle more realistic threats in practical applications. Moreover, understanding how the adversaries can exploit the interpretations derived from FedFM evaluation approaches to compromise the system is also crucial for the adoption of FedFM by mission-critical applications.
Efficiency and Accuracy Trade-offs: Existing FL evaluation techniques (e.g., shapely value, influence functions) incur high computation and communication costs if directly used to evaluate FedFM. This is especially challenging if resource-constrained devices are involved. Thus, research on trade-offs between efficiency and accuracy is important for FedFM evaluation for practical adoption.

### v2

@@ VI-D2 Contribution Evaluation for Centralized FMs
Automated Evaluation: Automated evaluation of FMs commonly adopts standard indicators or metrics and evaluation tools to assess model performance. These include accuracy, fairness, ROUGE [189] and BERTScore [190]. For instance, the BLEU score [191] has been leveraged to measure the similarity and quality between the text generated by the FM and the reference text by the machine translation task. Due to its simplicity and automatic computing, this evaluation metric is widely adopted by existing FM evaluation effort. In addition, with the deployment of FMs, more advanced automated evaluation approaches are emerging. For example, Lin et al. [192] proposed a unified multidimensional automated evaluation approach, LLM-EVAL, for domain conversations. PandaLM [193] trained an additional LLM to evaluate different models, which is suitable for reproducible and automated language model evaluation.
Human Evaluation: Human evaluation aims to assess the quality and accuracy of FM-generated results by human participation. Different from automated evaluation, human evaluation focuses more on specific application scenarios and can provide more comprehensive results. Usually, evaluators consist of experts, researchers or the target users. Recent human evaluation methods involve tasks including generation, summarization and analogical reasoning tasks. Bubeck et al. [194] conducted various human-crafted tests on GPT-4. It can be observed that GPT-4 performs close to or even exceeds human performance on various tasks. In practice, automated evaluation and human evaluation are considered for adoption on a case-by-case basis.
@@
During fine-tuning, the training data might not be equally important for a pre-trained FM. On the one hand, pre-trained FMs are prone to significant performance degradation with noisy data. This effect can be further amplified when noisy samples are highly influential to the model. On the other hand, specific knowledge embedded within some samples might have been extracted after several training rounds. Therefore, they can be ignored afterwards without affecting final fine-tuned model performance. Therefore, it is important to identify noisy samples and important samples in the fine-tuning data to improve training efficiency and model performance. Jain et al. [195] evaluate FMs from the data-based perspective by eliminating the need for laborious labeling of new data. In [196], an efficient SV-based data evaluation method has been proposed. It achieves this design goal through an efficient sampling-based method that aggregates SV values calculated from subsets for valuation of the entire training set, and a value transfer method that leverages value information extracted from a simple classifier trained by representations. DataInf [179] is a computationally efficient influence approximation method that is based on an easy-to-compute closed-form expression. It can be easily applied to FMs.
@@ VI-D3 Contribution Evaluation for FedFM
As for evaluating FedFM, so far, there have been few relevant works. The work [197] proposed a logic rule learning approach to select the optimal chain-of-thoughts prompts for improving the interpretability of federated prompt selection for multi-domain FMs. They cast this problem as a bi-level program, and solve it through variational expectation maximization. The work [198] proposed to evaluate FedFM by comparing the performance of the federated parameter-efficient fine-tuned model with traditional centralized fine-tuning methods. FedIT [199] leverages FL framework for FM instruction tuning and conducted studies on the widely used GPT-4 to exploit the heterogeneous and diverse sets of instructions. FedNLP [200] is a benchmark framework for evaluating FL methods on four common formulations of NLP tasks: text classification, sequence tagging, question answering and seq2seq generation. They proposed a universal interface between Transformer-based language models (e.g., BERT, BART) and FL methods under various non-IID partitioning strategies.
@@ VIII-B Communication Challenges facing FedFM
Increased Communication Overhead vs. Bandwidth Constraints: Bandwidth constraints present a critical challenge to FedFM [226]. The sheer size of FMs necessitates large transmission overhead, which is compounded by the need for frequent communication across distributed training devices (i.e., clients and servers). While strategies like model compression (using techniques like pruning, quantization, and knowledge distillation) aim to reduce the size of the models, and thus the network load [227], they only partially alleviate the bandwidth demand. Sparse communication techniques [228], which focus on sharing only essential model updates, help to some extent but do not entirely solve the problem. Asynchronous communication strategies [229] and edge computing [230] can reduce network congestion and data transmission volume. Yet, they cannot fully compensate for the inherent high transmission requirements of FedFM. Efficient coding schemes optimize data transmission [231], but the fundamental challenge of transmitting large volumes of data in a bandwidth-limited environment persists. Collectively, these approaches would make strides in addressing bandwidth issues, but they do not completely solve the challenge, highlighting the pressing need for innovative solutions in the field of FedFM [15].
@@
Privacy Issues: Privacy concerns in FedFM remain a significant challenge, despite the ongoing privacy protection research in traditional FL settings. While FL enhances privacy by sharing model updates rather than raw data, there are still vulnerabilities during interactions between FL participants [242]. For instance, adversaries can exploit these vulnerabilities to perform attacks such as membership inference or gradient leakage [242, 243, 244], aiming to extract local training data from devices. Existing countermeasures, such as homomorphic encryption (HE) and secure multi-party computation (SMC), while effective in enhancing FL privacy to some extent, fall short in fully addressing these malicious attacks. As detailed in [245], HE can protect against data leakage but is less effective against more sophisticated threats like membership inference and gradient leakage attacks. However, it cannot be efficiently applied on large-scale FMs. Despite these efforts in improving privacy protection, the evolving nature of threats and the complexity of FL systems, particularly in the context of FMs, means that privacy concerns remain a pressing and unresolved issue in FedFM. The development of more advanced and resilient privacy-preserving mechanisms continues to be a critical need in this field.
@@ VIII-C Promising Direction: Quantum Computing for FedFM
The integration of quantum computing with FedFM presents a transformative opportunity to enhance FedFM [246]. This integration could revolutionize the efficiency and capability of these models [247]. In this section, we envision how four promising quantum computing techniques (Table X): 1) quantum machine learning, 2) quantum optimization, 3) quantum security and 4) a combination of quantum entanglement and quantum aggregation hold promise to improve the efficiency, trustworthiness and incentives aspects of FedFM.
@@
Quantum Entanglement and Quantum Aggregation for FedFM: Quantum entanglement offers a novel approach to synchronizing updates in FedFM. This quantum phenomenon, where two or more particles become interconnected so that the state of one instantly influences the others, can be useful for FedFM in multiple ways. Theoretically, quantum entanglement might enable immediate synchronization of model updates across different nodes in a federated network [263]. Using entangled particles to represent model parameters allows changes in one node to be instantly reflected in all other entangled nodes, ensuring real-time synchronization of updates. Reaching consensus on model updates among distributed nodes is a significant challenge in FedFM. Quantum entanglement can achieve highly resource-efficient synchronization by reducing communication overhead and allowing for more effective error correction [264]. This can save bandwidth and computational resources, with quantum entanglement-assisted aggregation potentially correlating model updates from different nodes more efficiently. While still in the theoretical stage, the potential impact of quantum entanglement on FedFM could revolutionize the way model updates are synchronized, with highly enhanced efficiency, security and resource management.
Quantum aggregation in FedFM involves leveraging quantum computing principles for aggregating model updates from various nodes. This process aims to compute a global model update that effectively integrates local model changes from each node, potentially enhancing both efficiency and security. Key approaches in quantum aggregation include the follows. Quantum superposition [265] and quantum interference [266] involve using quantum superposition to represent multiple model updates on a single quantum state. This state can then be processed to compute a global model updating. Quantum interference, where quantum states interfere constructively or destructively, can facilitate efficient FedFM aggregation. Quantum summation algorithms can perform aggregation tasks more efficiently than classical algorithms [267]. Leveraging quantum cryptography techniques, quantum secure aggregation of model updates can preserve privacy while enabling effective aggregation. While detailed studies specific to FedFM and quantum aggregation are not yet available, the principles of quantum computing offer promising pathways for enhancing the efficiency and security of FedFM model aggregation.
@@ VIII-D Promising Direction: Evaluation of FedFM
Evaluation under Complex Threat Models: Existing FL evaluation methods are mostly based on the simple threat model of semi-honest participants. This renders them susceptible to scenarios in which the server or clients are colluding or malicious. It is imperative to relax this simplifying assumption to enable future FedFM evaluation methods to handle more realistic threats in practice. Moreover, understanding how the adversaries can exploit the interpretations derived from FedFM evaluation methods to compromise the system is also crucial for the adoption of FedFM by mission-critical applications.
Efficiency and Accuracy Trade-offs: Existing FL evaluation techniques (e.g., shapely value, influence functions) incur high computation and communication costs if directly used to evaluate FedFM. This is especially challenging if resource-constrained devices are involved. Therefore, research focusing on trade-off between efficiency and accuracy is very important for FedFM evaluation for practical adoption.